stone is thrown upwards with an initial velocity of 25 m/s at an angle of 30° to the ground. Show that the vertical component of the velocity at the start is 12.5 m/s up words
1
Expert's answer
2020-01-03T09:19:10-0500

The vertical component of the velocity


"v_y=v_0\sin\theta=25\times \sin 30^{\circ}=12.5\:\rm m/s"